# OmAgent: A Multi-modal Agent Framework for Complex Video Understanding with Task Divide-and-Conquer

### v2

@@ Abstract
Lu zhang1, Tiancheng Zhao1,2, Heting Ying1, Yibo Ma1, Kyusong Lee1,2 1Linker Technology Research, 2Binjiang Institute of Zhejiang University {Lu_zhang, Heting_Ying, Yibo_Ma}@hzlh.com   {tianchez, kyusongl}@zju-bj.com
@@ 1 Introduction
However, as LLMs have become more complex and powerful, there has been growing interest in applying them to more dynamic media, such as video content Lin et al. (2023a); Zhu et al. (2023); Zhang et al. (2023); Zhao et al. (2023); Tang et al. (2023). This interest arises from the potential to provide deeper and more nuanced interpretations of video data, similar to how humans understand and interact with moving images and sound. Currently, most video understanding models are limited to processing short videos, typically only a few minutes or even several seconds long. Despite these advancements, significant challenges remain, especially in handling long video inputs like 24-hour CCTV footage or full-length movies, which involve massive amounts of data and require substantial processing power.
@@ 2 Related Work
Analyzing and understanding video content using large-scale language models (LLMs) typically involves fine-tuning or pre-training methods. Pre-training strategies, such as supervised or contrastive learning, develop video LLMs, while instruction fine-tuning updates adapter parameters to enable video comprehension Tang et al. (2023). For example, LaViLa Zhao et al. (2023) enhances video subtitle generation through a cross-attention module and rewriting mechanism, improving coverage and diversity. Video-LLaMA Zhang et al. (2023) addresses spatio-temporal visual variations using separate video and audio encoders with an advanced audio-visual Q-former, significantly boosting video comprehension. Video-LLaVA Lin et al. (2023a) connects multimodal representations into a unified semantic space with LLMs, improving video understanding tasks.
@@
LLMs and Multimodal LLMs (MLLMs) applied to long video comprehension tasks utilize external systems to process extensive content. This involves analyzing visual elements, actions, scenes, and objects over time, aligning multimodal information with textual modalities, and leveraging the powerful text processing capabilities of MLLMs Tang et al. (2023). For instance, Vlog Kevin (2023) uses pre-trained models for different modalities to record and interpret visual and audio information, summarizing it into detailed text for MLLM comprehension. MM-REACT Yang et al. (2023) employs visual expert tools via internal prompts, enhancing MLLMs’ visual understanding. MM-VID Lin et al. (2023b) segments videos using ASR and Scene Detection tools, generating and integrating textual descriptions to complete Q&A tasks with MLLMs.
@@ 4.1 General problem-solving capabilities
The Mostly Basic Programming Problems (MBPP) benchmark compendium encompasses 976 elementary Python coding conundrums. These queries serve to test the system’s proficiency in strategizing solutions, its adroitness in selecting and invoking tools, as well as its resilience in rectifying errors.
FreshQA stands as an ongoing, perpetually refreshed repository of questions and answers from the open world, with its contents echoing the ceaseless evolution of reality, hence the ever-shifting sands of its responses. Therefore, FreshQA places a premium on the system’s capability to assimilate external information.
@@ 4.2 Long-form videos understanding capabilities
We created a benchmark with over 2000 question-answer pairs to evaluate OmAgent’s ability to understand, answer questions, and recall detailed information from long-form videos. This benchmark allows us to measure the capabilities of our intelligent agent and compare its performance against large language models.
@@
Publicly available long-form video understanding datasets are very scarce, similar to the FreshQA/MovieQA datasets which only contain videos at the minute level, and the questions are not complex enough, for example, "How does Talia die?" in MovieQA, which is a question that can be inferred from consecutive frames, cannot meet our needs. MoVQA is a long movie question and answer dataset that utilizes 100 well-known movies to create complex long video question-answer pairs, which align with our requirements, but it is not yet open-sourced. Therefore, we created the dataset ourselves. We collected some long videos familiar to annotators, selected the top 100 videos in terms of frequency, and created 20 questions for each video. These questions were first proofread by two different annotators and then revised by a third annotator. Videos include episodes and movies, varieties, documentaries, and vlogs. These types of videos exhibit significant differences in themes, filming and editing techniques, data density, scene lengths, and alignment of video audio and visuals, fully demonstrating the diversity of the data.
@@ 5.1 General problem-solving capabilities
OmAgent 57.14% 56.25% 23.53% 36.84% 45.45%
Table 2: Results of different types of videos on OmAgent and the other two control groups.
@@
Video2RAG 72.73% 4.76% 50.17% 23.36%
OmAgent 81.82% 19.05% 72.74% 57.21%
Table 3: Results of different types of queries on OmAgent and the other two control groups.
@@ 5.2 Long-form videos understanding capbilities
Table 2 compares the scores of two control groups and OmAgent across different types of long-form video understanding. OmAgent achieved the highest scores. Vlogs, variety, and documentaries contain extensive narration, so the STT data and the resulting scene captions encompass most relevant information of the video. Therefore, the performance difference between OmAgent and the other two methods in these categories is not as significant as in episodes and movies. In episodes and movies, scenes change frequently, complex queries involve cross-scene information, and STT data might span scene transitions. Compared to frames with STT, Video2RAG retrieves data related to the query, reducing data redundancy, and thus has higher scores than frames with STT. However, since it only retrieves relevant information from a vector database, complex questions such as "Are there any scene changes between 03:58 and 04:02, and what is their connection?" are not achieved. On the other hand, OmAgent’s DnC Loop breaks down complex questions into several sub-questions, including "Extract frames between 03:58 and 04:02," "Analyze the extracted frames to identify any scene changes," and "Determine the connection between the scenes based on the identified changes." By leveraging the rewinder capability, it pinpoints the relevant segments for rewatching, thereby arriving at the correct answer.
@@
Furthermore, we conducted a detailed analysis of different question types. Table 3 provides a comparison of OmAgent and two control groups in terms of reasoning, event localization, information summary, and external knowledge. The results show that OmAgent achieves the highest scores in all four types of questions. Through its rewinder capability, OmAgent can extract more detailed video information and accurately locate timestamps, which leads to significant improvements in reasoning, event localization, and information summarization tasks compared to frames with STT and Video2RAG. External knowledge task has stricter requirements for information retrieval. Although GPT can answer some questions through its own capabilities and scene information, OmAgent achieves higher scores by utilizing various external tools (such as facial recognition, web search, etc.) to obtain more accurate relevant information. Notably, in the question type of information summary, Video2RAG scored lower than frames with STT. Analysis reveal that video2RAG’s information source came from scene captions generated based on STT and images within the scene, which had a certain probability of omissions compared to the original STT, thus affecting the scores. Although OmAgent performed the best across all question types, it still scores relatively low in the event localization task. Detailed study and analysis reveal that LLM tends to directly use timestamps or time spans within the scene, leading to a lack of precision in answers. OmAgent’s DnC loop and rewinder capabilities can mitigate this issue but cannot completely resolve it.

### v3

@@ Abstract
Lu Zhang1, Tiancheng Zhao1,2, Heting Ying1, Yibo Ma1, Kyusong Lee1,2 1Om AI Research, 2Binjiang Institute of Zhejiang University {zhang_lu, ying_heting, ma_yibo}@hzlh.com   {tianchez, kyusongl}@zju-bj.com
@@ 1 Introduction
However, as LLMs have become more complex and powerful, there has been growing interest in applying them to more dynamic media, such as video content Lin et al. (2023a); Zhu et al. (2023); Zhang et al. (2023b); Zhao et al. (2023); Tang et al. (2023). This interest arises from the potential to provide deeper and more nuanced interpretations of video data, similar to how humans understand and interact with moving images and sound. Currently, most video understanding models are limited to processing short videos, typically only a few minutes or even several seconds long. Despite these advancements, significant challenges remain, especially in handling long video inputs like 24-hour CCTV footage or full-length movies, which involve massive amounts of data and require substantial processing power.
@@ 2 Related Work
Analyzing and understanding video content using large-scale language models (LLMs) typically involves fine-tuning or pre-training methods. Pre-training strategies, such as supervised or contrastive learning, develop video LLMs, while instruction fine-tuning updates adapter parameters to enable video comprehension Tang et al. (2023). For example, LaViLa Zhao et al. (2023) enhances video subtitle generation through a cross-attention module and rewriting mechanism, improving coverage and diversity. Video-LLaMA Zhang et al. (2023b) addresses spatio-temporal visual variations using separate video and audio encoders with an advanced audio-visual Q-former, significantly boosting video comprehension. Video-LLaVA Lin et al. (2023a) connects multimodal representations into a unified semantic space with LLMs, improving video understanding tasks.
@@
LLMs and Multimodal LLMs (MLLMs) applied to long video comprehension tasks utilize external systems to process extensive content. This involves analyzing visual elements, actions, scenes, and objects over time, aligning multimodal information with textual modalities, and leveraging the powerful text processing capabilities of MLLMs Tang et al. (2023). For instance, Vlog Kevin (2023) uses pre-trained models for different modalities to record and interpret visual and audio information, summarizing it into detailed text for MLLM comprehension. MM-REACT Yang et al. (2023) employs visual expert tools via internal prompts, enhancing MLLMs’ visual understanding. MM-VID Lin et al. (2023b) segments videos using ASR and Scene Detection tools, generating and integrating textual descriptions to complete Q&A tasks with MLLMs. LLoVi Zhang et al. (2023a) uses a process that generates a summary based on subtitles and questions, then uses the summary for question-answering. VideoTree Wang et al. (2024b) employs a three-step process to understand long videos, clustering video frames, calculating relevance, and performing depth expansion for question-answering. VideoAgent Fan et al. (2024) preprocesses videos to generate captions and object data, and uses an agent with pre-provided tools to obtain answers.
@@ 4.1 General problem-solving capabilities
The Mostly Basic Programming Problems (MBPP) benchmark includes 976 elementary Python coding tasks. These problems are designed to evaluate the system’s proficiency to plan solutions, select and invoke the right tools, and fix errors effectively.
FreshQA is a continuously updated collection of real-world questions and answers, reflecting the constantly changing nature of reality. As a result, FreshQA focuses on the system’s ability to learn and integrate new information from external sources.
@@ 4.2 Long-form videos understanding capabilities
We created a benchmark with over 2000 question-answer pairs to evaluate OmAgent’s ability to understand, answer questions, and recall detailed information from long-form videos. We aimed to assess the general understanding of ultra-long video content and the ability to recall specific details. Our benchmark was designed to be logically coherent and narrative-rich, with video segments up to an hour long. This benchmark allows us to measure the capabilities of our intelligent agent and compare its performance against large language models.
@@
Publicly available long-form video understanding datasets are very scarce, similar to the FreshQA/MovieQA datasets which only contain videos at the minute level, and the questions are not complex enough, for example, "How does Talia die?" in MovieQA, which is a question that can be inferred from consecutive frames, cannot meet our needs. SOK-Bench Wang et al. (2024a) addresses scenarios slightly different from ours, as it demands the testing program to integrate situated and general knowledge to answer questions. MoVQA is a long movie question and answer dataset that utilizes 100 well-known movies to create complex long video question-answer pairs, which align with our requirements, but it is not yet open-sourced. Therefore, we created the dataset ourselves. We collected some long videos familiar to annotators, selected the top 100 videos in terms of frequency, and created 20 questions for each video. These questions were first proofread by two different annotators and then revised by a third annotator. Videos include episodes and movies, varieties, documentaries, and vlogs. These types of videos exhibit significant differences in themes, filming and editing techniques, data density, scene lengths, and alignment of video audio and visuals, fully demonstrating the diversity of the data.
@@ 5.1 General problem-solving capabilities
Frames with STT 42.85% 29.41% 17.64% 31.58% 28.57%
VideoAgent 41.72% 23.53% 11.76% 26.32% 23.38%
VideoTree 34.52% 31.48% 21.27% 27.35% 26.76%
LLoVi 28.57% 24.16% 17.65% 21.05% 23.63%
Table 2: Results of different types of videos on OmAgent and the other five baselines.
@@
VideoAgent 64.66% 2.25% 45.45% 23.78%
VideoTree 35.30% 18.62% 47.27% 29.57%
LLoVi 27.27% 11.90% 45.46% 24.57%
Table 3: Results of different types of queries on OmAgent and the other five baselines.
@@ 5.2 Long-form videos understanding capbilities
Table 2 compares the scores of five baselines and OmAgent across different types of long-form video understanding. OmAgent achieved the highest scores. Vlogs, variety, and documentaries contain extensive narration, so the STT data and the resulting scene captions encompass most relevant information of the video. Therefore, the performance difference between OmAgent and the other two methods in these categories is not as significant as in episodes and movies. In episodes and movies, scenes change frequently, complex queries involve cross-scene information, and STT data might span scene transitions. Compared to frames with STT, Video2RAG retrieves data related to the query, reducing data redundancy, and thus has higher scores than frames with STT. However, since it only retrieves relevant information from a vector database, complex questions such as "Are there any scene changes between 03:58 and 04:02, and what is their connection?" are not achieved. On the other hand, OmAgent’s DnC Loop breaks down complex questions into several sub-questions, including "Extract frames between 03:58 and 04:02," "Analyze the extracted frames to identify any scene changes," and "Determine the connection between the scenes based on the identified changes." By leveraging the rewinder capability, it pinpoints the relevant segments for rewatching, thereby arriving at the correct answer.
@@
Furthermore, we conducted a detailed analysis of different question types. Table 3 provides a comparison of OmAgent and five baselines in terms of reasoning, event localization, information summary, and external knowledge. The results show that OmAgent achieves the highest scores in all four types of questions. Through its rewinder capability, OmAgent can extract more detailed video information and accurately locate timestamps, which leads to significant improvements in reasoning, event localization, and information summarization tasks compared to frames with STT and Video2RAG. External knowledge task has stricter requirements for information retrieval. Although GPT can answer some questions through its own capabilities and scene information, OmAgent achieves higher scores by utilizing various external tools (such as facial recognition, web search, etc.) to obtain more accurate relevant information. Notably, in the question type of information summary, Video2RAG scored lower than frames with STT. Analysis reveal that video2RAG’s information source came from scene captions generated based on STT and images within the scene, which had a certain probability of omissions compared to the original STT, thus affecting the scores. The higher frame extraction frequency of VideoAgent, VideoTree and LLoVi allowed them to capture more visual information, but their lack of audio processing led to suboptimal performance in many scenarios. VideoAgent’s lack of DnC and rewinder capabilities limited its effectiveness in query types requiring detailed information comparing with OmAgent. Although OmAgent performed the best across all question types, it still scores relatively low in the event localization task. Detailed study and analysis reveal that LLM tends to directly use timestamps or time spans within the scene, leading to a lack of precision in answers. OmAgent’s DnC loop and rewinder capabilities can mitigate this issue but cannot completely resolve it.
@@ 7 Limitations
8 Acknowledgements
This research is partially supported by National Key R&D Program of China under grant No. 2022YFF0902600